OpenStudy (rizags):

how do i properly express the domain of the function f(x) = 2/sqrt(x^2-5) in set builder notation

OpenStudy (rizags):

\[f(x) = \frac{ 2 }{ \sqrt{x^{2}-5} }\]

OpenStudy (rizags):

i know the domain is all reals except \[\sqrt{5} and -\sqrt{5}\]

OpenStudy (rizags):

but how do I express it properly?

jimthompson5910 (jim_thompson5910):

keep in mind that you cannot take the square root of a negative number

OpenStudy (rizags):

but wait if x is being squared, then it doesnt matter if x is negative, becuase -sqrt(5) squared is 5 right?

jimthompson5910 (jim_thompson5910):

yes but x^2 - 5 is negative for some interval

jimthompson5910 (jim_thompson5910):

you have to make sure x^2 - 5 is NOT zero and also NOT negative

OpenStudy (rizags):

the original question had x squared Plus 5, sorry

OpenStudy (rizags):

thats what i meant, i will redrwaw it

OpenStudy (rizags):

\[f(x) = \frac{ 2 }{ \sqrt{x^{2}+5} }\]

OpenStudy (anonymous):

\[x = \left\{ x \in \mathbb{R} | \left| x \right| > \sqrt{5} \right\}\]

OpenStudy (anonymous):

Oh, just seen what you said about it being +5, ignore the above.

jimthompson5910 (jim_thompson5910):

you are correct though tom

jimthompson5910 (jim_thompson5910):

that's the same as saying x < -sqrt(5) or x > sqrt(5)

OpenStudy (rizags):

ok. Becuase in my test, i wrote the answer as the domain of \[D _{x}: \mathbb{R}/-\sqrt{5},\sqrt{5}\] and was marked wrong

OpenStudy (rizags):

i dont know why

jimthompson5910 (jim_thompson5910):

yeah because you don't just kick out sqrt(5), -sqrt(5)

jimthompson5910 (jim_thompson5910):

you have to kick out everything in between because those values make x^2 - 5 > 0 false

OpenStudy (anonymous):

Thanks @jim_thompson5910, but surely the domain is all real numbers excluding +/-sqrt(5) now that it's x^2+5 in the denominator?

OpenStudy (rizags):

wait, suppose i was to kick out only the values of -sqrt(5) and sqrt(5) for some function g, how would i write that domain?

OpenStudy (rizags):

jim, i am no longer using the x^2 - 5, now its the x^2 + 5

jimthompson5910 (jim_thompson5910):

oh it's been changed to +5, I see

OpenStudy (anonymous):

\[\left\{ x \in \mathbb{R} | \left| x \right| \neq \sqrt{5}\right\}\]

jimthompson5910 (jim_thompson5910):

so x^2 + 5 is ALWAYS positive x^2 is nonnegative, adding a positive bumps it into an all positive region

jimthompson5910 (jim_thompson5910):

x^2+5 being always positive means x^2 + 5 = 0 has no solutions

OpenStudy (anonymous):

@Rizags, your thinking was correct but I think the problem is that you haven't written your answer in set builder notation. Maybe Jim can confirm this?

jimthompson5910 (jim_thompson5910):

the domain is the set of all real numbers there is no real number x that causes a division by zero error since x^2 + 5 = 0 has no real solutions we also don't have to worry about taking the square root of a negative number

OpenStudy (rizags):

it isnt about set builder, the teacher said that my answer represented a disjunction of the values, but the answer should have been a conjuction, as in x cannot equal sqrt(5) AND -sqrt(5). Just not sure how to write that.

OpenStudy (rizags):

and she disregarded the fact that x^2 + 5 has no real solutions

OpenStudy (anonymous):

Oops, I'm back on x^2-5 with my thinking, lol. Thanks for pointing that out, Jim.

jimthompson5910 (jim_thompson5910):

the graph visually confirms it

OpenStudy (rizags):

Essentially, what I am asking is, for any function g, with domain of all values excluding sqrt(5) and -sqrt(5), how would i write the domain in set builder without using absolute value of x, as tom did above?

OpenStudy (anonymous):

@Rizags I was wrong. What Jim is saying is that it's impossible to find a real number that doesn't work in your equation (try putting sqrt(5) in and you'll get sqrt(10)/5 as your output) - hence your domain is R, the real numbers.

OpenStudy (rizags):

And one more thing, is this the proper way to represent the limit of the function \[x ^{2} - 4x + 7\] \[\lim_{h \rightarrow 0}\frac{ f(x+h) - f(x) }{ h } = 2x+4 \]

OpenStudy (rizags):

how would i express the difference quotient? is this coorect?

OpenStudy (rizags):

is that equation i wrote correct?

jimthompson5910 (jim_thompson5910):

f(x) = x^2 - 4x + 7 you need to find f(x+h)

jimthompson5910 (jim_thompson5910):

2x+4 is close, but incorrect

OpenStudy (rizags):

oh, its minus 4

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (rizags):

but is the equation format correct, because my teacher is very strict about that

OpenStudy (rizags):

would i write, when F(x) = blank, then the limit equation?

jimthompson5910 (jim_thompson5910):

you have a true statement if you write \[\Large \lim_{h \rightarrow 0}\frac{ f(x+h) - f(x) }{ h } = 2x-4\]

OpenStudy (rizags):

but, i have to identify f(x) first, right?

jimthompson5910 (jim_thompson5910):

of course there are multiple steps in between, but it's still correct

jimthompson5910 (jim_thompson5910):

f(x) = x^2 - 4x + 7

OpenStudy (rizags):

but do i identify that in my formal answer?

jimthompson5910 (jim_thompson5910):

you find f(x+h) and then simplify and then find the limit

OpenStudy (rizags):

or is it implied?

jimthompson5910 (jim_thompson5910):

I'm not sure what your teacher wants ultimately

jimthompson5910 (jim_thompson5910):

but s/he probably wants the steps to get from A to Z

OpenStudy (rizags):

i know how to do it, its just i need a very clear representation of the answer. Would this be clear? \[When f(x) = x^{2} - 4x + 7,\] \[\lim_{h \rightarrow 0}\frac{ f(x+h)-f(x) }{ h } = 2x - 4\]

jimthompson5910 (jim_thompson5910):

yes that is clear and correct

OpenStudy (anonymous):

I can't imagine she will have any problems with that, it looks good to me.

OpenStudy (rizags):

ok, thanks a lot for the help I appreciate it

jimthompson5910 (jim_thompson5910):

np